# mystylerightline=true, innerleftmargin=10, innerrightmargin=10, outerlinewidth=3pt, topline=false, rightline=false, bottomline=false, skipabove=skipbelow=false

### v1

@@ 7. Related Work
A) Pre-processing Techniques exploit the space of input data to mitigate fairness defects (Galhotra et al., 2022; Feldman et al., 2015; Kamiran and Calders, 2012; Zemel et al., 2013). Reweighting (Kamiran and Calders, 2012) augments the dataset where the datapoints in each group-label combination are weighted differently to ensure fairness. The reweighting process minimizes the discrepancy between the observed and expected probabilities of the favorable outcome occurring with the discriminated sensitive attribute. FairMask (Peng et al., 2022) uses the fact that protected attributes can be inferred by some combinations of other non-protected attributes (the “proxy” problem) to reduce the influence of protected attributes in the inference stage. Fair-SMOTE (Chakraborty et al., 2021a) used under-sampling and over-sampling techniques (Chawla et al., 2002) to balance data based on class and sensitive attributes. MAAT (Chen et al., 2022) proposes an ensemble training approach to mitigating unfairness. They propose a framework to train fairness and performance ML models and combine their outputs to make predictions. The performance model is identical to traditional ML algorithms and remains unchanged. For the fairness model, they design a debugging strategy based on prior work (Wick et al., 2019; Chakraborty et al., 2021b) that attribute bias in training data to selection bias and label bias. LTDD (Li et al., 2022) hypothesizes that biased features in training data contributes to unfairness. To obtain unbiased features, they develop a linear regression based algorithm that measures the association between non-sensitive and sensitive features in the dataset. The non-sensitive features are then modified to remove the biased parts to become independent of the sensitive features. After debiasing, the sensitive features are dropped from the dataset.
@@
B) In-processing techniques aim to improve fairness during training. One method, known as adversarial debiasing (Zhang et al., 2018), employs adversarial learning to develop a classifier that not only improves prediction accuracy but also hinders the ability of adversaries to determine sensitive attributes from ML predictions. Another approach, the prejudice remover (Kamiran et al., 2012), incorporates a fairness-centric adjustment into the learning objective (loss function), aiming to balance both accuracy and fairness. These strategies necessitate alterations in either the loss function or the parameters of the ML models. Seldonian (Thomas et al., 2019; Metevier et al., 2019; Hoag et al., 2023) presented a technique that allows users to directly specify (arbitrary) undesirable behaviors as constraints and enforce them during training. Fairway (Chakraborty et al., 2020) combines pre-processing and in-processing mitigation techniques to improve fairness. Parfait-ML (Tizpaz-Niari et al., 2022) is a gray-box evolutionary search algorithm that explores the ML hyperparameters to find configurations that minimize fairness while maintaining an acceptable accuracy. While mitigating unfairness during training might be less intrusive than changing the training datasets; they often require changing training algorithms that might be only effective for a particular algorithm and fairness definition. In addition, the training process is often expensive and might not be feasible for existing systems.
@@
C) Post-processing techniques aim to modify the prediction outcomes of ML models to reduce discrimination (Hardt et al., 2016; Kamiran et al., 2012; Pleiss et al., 2017; Tao et al., 2022). Equalized Odds Processing (EOP) (Hardt et al., 2016) is a post-hoc correction mechanism that modifies the output of a biased binary classifier. EOP solves an optimization problem using a linear program that utilizes the protected attribute, predictions, and true labels from a dataset to create an unbiased predictor. NeuFair, however, takes a fundamentally different approach to post-processing mitigation. First, NeuFair does not require training auxiliary models to mitigate unfairness. Second, NeuFair does not rely on randomization for fairness and generates a deterministic DNN classifier. Predictions during inference time are deterministic after the neurons are dropped from the DNN. Lastly, Woodworth et al. (2017) shows that EOP requires the biased classifier to be Bayes optimal, which is practically impossible to learn from finite samples of data. Zhang and Sun (Zhang and Sun, 2022) adaptively intervened in the input data features and DNN internal neurons to improve fairness. Care (Sun et al., 2022) presented Particle Swarm Optimisation (PSO) to repair DNNs. Faire (Li et al., 2023) proposes a method to repair unfair DNNs by altering the activations of neurons. Inspired by program repair techniques, Faire first groups neurons representing protected and non-protected features in the dataset into two categories. To create the two categories, Faire creates a clone of the unfair model and retrains it to predict the protected attribute. By comparing the activations of the corresponding neurons in both networks, Faire determines which neurons represent the protected and non-protected. Finally, the protected neuron activations are penalized, and the non-protected neuron activations are promoted. Compared to Faire, NeuFair does not retrain the model and does not separately use the protected attribute to identify unfair neurons. We find an unfair subset of neurons given the entire feature vector, thereby accounting for the causal influence of the protected features on the non-protected features in the training dataset. Our analysis considers a group of neurons together rather than identifying the contributions of individual neurons. FairNeuron (Gao et al., 2022) proposes an algorithm to selectively retrain neurons in a neural network to mitigate unfairness. On a high level, the algorithm identifies input-output neuron paths corresponding to each sample in the training dataset that induces bias in the prediction. Then, they split the dataset into biased and unbiased subsets corresponding to the neuron activations in the paths. Finally, the algorithm re-trains the model with both training data splits by enabling random dropouts on the neurons in the biased paths and disabling any dropouts for neurons in the un-biased paths. While FairNeuron utilizes dropouts to improve fairness, it fundamentally differs from NeuFair. First, we do not require access to the large training dataset, and we do not retrain the models. Second, NeuFair algorithms utilize deterministic dropout during inference time to mitigate unfairness with a flexible budget on the loss in F1 score, i.e., we identify a desirable subset of neurons to drop that has a minimal impact on F1 score but the maximal impact on fairness, rather than dropping neurons randomly. Lastly, since we do not perform re-training, NeuFair does not update the weights of any neuron in the network.
@@ 8. Conclusion
In this paper, we tackle the problem of mitigating unfairness in pre-trained deep neural networks using the dropout method. We showed that the neural dropout problem over the DNN models is computationally hard and presented NeuFair, a family of randomized algorithms to efficiently and effectively improve the fairness of DNNs. Our experiments showed that NeuFair can identify an ideal subset of neurons to drop that disparately contribute to unfairness (leading to up to 69% fairness improvement) and outperform a state-of-the-art post-processing bias mitigator.
For future work, there are a few exciting directions. First, instead of choosing the best state found during the search, we can analyze the top $k$ states found during the run. Such a strategy could reduce the difference in improvement observed between the validation and test sets. We can extend to multi-valued protected attributes (age groups) or optimize for more than one protected attribute at a time. Second, we plan to leverage dropouts in the fine-tuning process to understand the effects of randomized algorithms on mitigating the unfairness and toxicity of large language models. Lastly, alternate strategies like genetic algorithms can be considered to solve the combinatorial optimization problem of neuron dropout. We use randomized algorithms for their simplicity in efficiently solving combinatorial optimization problems. The primary difference between SA/RW and genetic algorithms is that genetic algorithms create a population of candidate solutions at each step instead of a single solution. Genetic algorithms can be promising as they enable us to explore more states of the search space simultaneously.
Data Availability. Our open-source tool NeuFair with all experimental subjects are publicly accessible: Link.

### v2

@@ 7. Related Work
A) Pre-processing Techniques exploit the space of input data to mitigate fairness defects (Galhotra et al., 2022; Feldman et al., 2015; Kamiran and Calders, 2012; Zemel et al., 2013). Reweighting (Kamiran and Calders, 2012) augments the dataset where the data points in each group-label combination are weighted differently to ensure fairness. The reweighting process minimizes the discrepancy between the observed and expected probabilities of the favorable outcome occurring with the sensitive attribute. FairMask (Peng et al., 2022) uses the fact that protected attributes can be inferred by some combinations of other non-protected attributes (the “proxy” problem) to reduce the influence of protected attributes in the inference stage. Fair-SMOTE (Chakraborty et al., 2021a) used under-sampling and over-sampling techniques (Chawla et al., 2002) to balance data based on class and sensitive attributes. MAAT (Chen et al., 2022) proposes an ensemble training approach to mitigating unfairness. They propose a framework to train fairness and performance ML models and combine their outputs to make predictions. The performance model is identical to traditional ML algorithms and remains unchanged. For the fairness model, they design a debugging strategy based on prior work (Wick et al., 2019; Chakraborty et al., 2021b) that attribute bias in training data to selection bias and label bias. LTDD (Li et al., 2022) hypothesizes that biased features in training data contribute to unfairness. They develop a linear regression-based algorithm that measures the association between non-sensitive and sensitive features in the dataset to obtain unbiased features. The non-sensitive features are then modified to remove the biased parts to become independent of the sensitive features. The sensitive features are dropped from the training dataset after debiasing.
@@
B) In-processing techniques aim to improve fairness during training. Adversarial debiasing (Zhang et al., 2018), employs adversarial learning to develop a classifier that not only improves prediction accuracy but also hinders the ability of adversaries to determine sensitive attributes from ML predictions. Another approach, the prejudice remover (Kamiran et al., 2012), incorporates a fairness-centric adjustment into the loss function to balance accuracy and fairness. These strategies necessitate alterations in either the loss function or the model parameters. Seldonian (Thomas et al., 2019; Metevier et al., 2019; Hoag et al., 2023) presented a technique that allows users to directly specify (arbitrary) undesirable behaviors as constraints and enforce them during training. Fairway (Chakraborty et al., 2020) combines pre-processing and in-processing mitigation techniques to improve fairness. Parfait-ML (Tizpaz-Niari et al., 2022) is a gray-box evolutionary search algorithm that explores the ML hyperparameters to find configurations that minimize fairness while maintaining an acceptable accuracy. While mitigating unfairness during training might be less intrusive than changing the training datasets, it often requires changing the training algorithms. This might only be effective for a particular algorithm and fairness definition. Additionally, the training process is often expensive and might not be feasible for existing systems.
@@
C) Post-processing techniques aim to modify the prediction outcomes of ML models to reduce discrimination (Hardt et al., 2016; Kamiran et al., 2012; Pleiss et al., 2017; Tao et al., 2022). Equalized Odds Processing (EOP) (Hardt et al., 2016) is a post hoc correction mechanism that modifies the output of a biased binary classifier. EOP solves an optimization problem using a linear program that utilizes the protected attribute, predictions, and true labels from a dataset to create an unbiased predictor. NeuFair, however, takes a fundamentally different approach to post-processing mitigation. First, NeuFair does not require training auxiliary models to mitigate unfairness. Second, NeuFair does not rely on randomization for fairness and generates a deterministic DNN classifier. Predictions during inference time are deterministic after the neurons are dropped from the DNN. Lastly, Woodworth et al. (2017) shows that EOP requires the biased classifier to be Bayes optimal, which is practically impossible to learn from finite samples of data. Zhang and Sun (Zhang and Sun, 2022) adaptively intervened in the input data features and DNN internal neurons to improve fairness. Care (Sun et al., 2022) presented a Particle Swarm Optimization (PSO) algorithm to repair DNNs. Faire (Li et al., 2023) proposes a method to repair unfair DNNs by altering the activations of neurons. Inspired by program repair techniques, Faire first groups neurons representing protected and non-protected features in the dataset into two categories. To create the two categories, Faire creates a clone of the unfair model and retrains it to predict the protected attribute. By comparing the activations of the corresponding neurons in both networks, Faire determines which neurons represent the protected and non-protected. Finally, the protected neuron activations are penalized, and the non-protected neuron activations are promoted. NeuFair, however, does not retrain the model and does not separately use the protected attribute to identify unfair neurons. We find an unfair subset of neurons given the entire feature vector, thereby accounting for the causal influence of the protected features on the non-protected features in the training dataset. Our analysis considers a group of neurons together rather than identifying the contributions of individual neurons. FairNeuron (Gao et al., 2022) proposes an algorithm to mitigate unfairness by selectively retraining neurons in a DNN. On a high level, the algorithm identifies input-output neuron paths corresponding to each sample in the training dataset that induces bias in the prediction. Then, the dataset is split into biased and unbiased subsets corresponding to the neuron activations in the paths. Finally, the algorithm retrains the model with both training data splits by enabling random dropouts in the biased paths and disabling any dropouts in the unbiased paths. While FairNeuron utilizes dropouts to improve fairness, it fundamentally differs from NeuFair. First, we do not require access to the large training dataset, and we do not retrain the models. Second, NeuFair algorithms utilize deterministic dropout during inference time to identify a desirable subset of neurons to drop that has a minimal impact on F1 score but the maximal impact on fairness, rather than dropping neurons randomly. Lastly, NeuFair does not retrain the DNN to update weights.
@@ 8. Conclusion
In this paper, we tackle the problem of mitigating unfairness in pre-trained DNNs using the dropout method. We showed that the neural dropout problem over the DNN models is computationally hard and presented NeuFair, a family of randomized algorithms to efficiently and effectively improve the fairness of DNNs. Our experiments showed that NeuFair can identify an ideal subset of neurons to drop that disparately contribute to unfairness (leading to up to 69% fairness improvement) and outperform a state-of-the-art post-processing bias mitigator.
For future work, there are a few exciting directions. First, we can analyze the top $k$ states found during the run instead of choosing the best state. Such a strategy could reduce the difference in improvement observed between the validation and test sets. We can extend to multi-valued protected attributes (age groups) or optimize for more than one protected attribute at a time. Second, we plan to leverage dropouts in the fine-tuning process to understand the effects of randomized algorithms on mitigating the unfairness and toxicity of large language models. Lastly, alternate strategies like genetic algorithms can be used to solve the combinatorial optimization problem of neuron dropout. We use randomized algorithms for their simplicity in efficiently solving combinatorial optimization problems. The primary difference between SA/RW and genetic algorithms is that genetic algorithms create a population of candidate solutions at each step instead of a single solution. This approach can be promising as it enables us to explore more states of the search space simultaneously.
Data Availability
Our open-source tool NeuFair with all experimental subjects are publicly accessible: https://github.com/vdasu/neufair.
Acknowledgements.
This material is based upon work supported by the National Science Foundation under Grant No. CNS-2230061.